# A Comprehensive Survey on Uncertainty Quantification for Deep Learning

### v4

@@ 5.1.2. Ensemble models
Ensemble models combine multiple neural network models in the prediction process. The ensemble predictions form an output distribution. The prediction variability of the ensemble models can be a measure of model uncertainty (e.g., a larger variance implies larger uncertainty). To capture the model uncertainty arising from different aspects, several strategies for constructing ensembles can be adopted, which are summarized into three major categories: the first kind is through bootstrapping (Lakshminarayanan et al., 2017). This is a strategy for random sampling from the original dataset with replacement. An ensemble of neural network models is constructed and each model is trained on different bootstrapped samples. After training, the inference is done through the aggregation of the ensembles, and uncertainty is obtained from the prediction variance (regression) or average entropy (classification). The second strategy is to construct different neural network architectures (number of layers, hidden neurons, type of activation functions.) (Mallick et al., 2022). This strategy can account for the uncertainty from model misspecification. Other strategies include different initialization of parameters along with a random shuffle of the datasets. This is better than the bootstrap strategy since more samples are utilized for each model. The third type is hyperensemble (Wenzel et al., 2020). This approach constructs ensembles with different hyper-parameters, such as learning rate, optimization strategy, and training strategy.
@@
5.1.3. Sample density-aware neural network
Because of the theoretical soundness of the BNN model and the simplicity of deep ensemble models, they are the popular methods for modern DNN uncertainty quantification. However, the computational challenge with BNN and ensemble models make it infeasible for real-world applications. In addition, the approaches can address the model uncertainty associated with the parameter, model architecture, and training process stochasticity, but cannot generalize to model uncertainty coming from low sample density, which means, the samples lie far away from the support of the training sets may show over-confident results. In this regard, many approaches have been motivated to develop sample density-aware neural networks to capture the model uncertainty due to low training sample density.
@@ 5.1.3. Sample density-aware neural network
Then the prediction on new test input sample $\boldsymbol{x}_{*}$ is through computing the posterior distribution conditioned on the training data $\mathcal{D}_{\text{tr}}$, given by Eq. 11. GP inference gives lower uncertainty if the test samples are around the region where training samples are abundant (higher sample density as in the middle part of Fig. 6), otherwise resulting in higher uncertainty (boundary part of Fig. 6).
(11) $\footnotesize p(y_{*}|x_{*},\mathcal{D}_{\text{train}},\boldsymbol{\theta})=\mathcal{N}(y|\mathbf{K}_{\boldsymbol{x}}^{T}\mathbf{K}_{n}^{-1}\boldsymbol{y},K_{x*}-\mathbf{K}_{\boldsymbol{x}}^{T}\mathbf{K}_{n}\mathbf{K}_{\boldsymbol{x}}),$
@@
Sparse Gaussian process: Though possessing the intriguing property of inherent uncertainty estimation, GP is prohibitive for large datasets because the inversion of covariance matrix requires $\mathcal{O}(n^{3})$ complexity ($n$ is the total number of training samples), which takes high computation and storage cost for training and inference on large datasets. In this regard, many methods (Snelson and Ghahramani, 2005; Titsias, 2009)attempt to make a sparse approximation to the full GP to bring down the computation to $\mathcal{O}(m^{2}n)$ ($m$ is the number of inducing valriables and $m\ll n$. The inducing variables can be anywhere in the input domain, not constrained to be a subset of the training data, which is denoted as input-output pairs $\{\hat{\boldsymbol{x}}_{i},\hat{y}_{i}\}_{i=1}^{m}$. Then the inversion of the original covariance matrix $\mathbf{K}_{n}$ can be replaced with a low-rank approximation from the inducing variables, which only requires the inversion of $m\times m$ matrix $\mathbf{K}_{m}$. Then the question becomes how to select $m$ best-inducing variables to be representative of the training dataset. Common approaches assume that the best representative inducing variables are those that maximize the likelihood (ML) of the training data (Snelson and Ghahramani, 2005). Then the location of inducing variables and hyper-parameters of GP are optimized simultaneously through ML. The training data likelihood can be obtained by marginalizing out the inducing variables on the joint distribution of training data and inducing variables. Another approach variation sparse GP formulates a variation lower bound of the exact likelihood by treating the locations of inducing variables as variational parameters (Titsias, 2009) for optimization. To further reduce the computation for more scalable inference on a large dataset, the paper (Wilson and Nickisch, 2015) proposes a kernel interpolation method for scalable structured GP method by exploiting the structure of the covariance by imposing grid constraint on the inducing variables. In this way, the kernel matrix $\mathbf{K}_{m}$ admits the Kronecker structure and is much easier for computing the inversion.
@@
Deep Gaussian process: Besides the computational challenge, another limitation of GP is the joint Gaussian distribution assumption on the target variables limits the model’s capability in capturing diverse relationships among instances in large datasets. Additionally, GP relies heavily on the kernel function to compute the similarity between samples by transforming input features into the high-dimensional manifold. However, for high-dimension structured data, it is challenging to construct appropriate kernel functions to extract hierarchical features for computing similarity between samples. To address these limitations, two lines of researches area have been proposed: deep kernel learning and Deep Gaussian process.
@@
Deep kernel learning (Wilson et al., 2016) aims to combine the structured feature learning capability of DNN with the GP to learn more flexible representations. The motivation is that DNN can automatically discover meaningful representations from high-dimensional data, which could alleviate the fixed kernel limitations of GP and improve its expressiveness. Specifically, the deep kernel learning transforms the kernel $\mathbf{K}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ to $\mathbf{K}_{\boldsymbol{\theta}}(g(\boldsymbol{x}_{i};\boldsymbol{w}),g(\boldsymbol{x}_{j},\boldsymbol{w}))$, where $g(\cdot;\boldsymbol{w})$ is the neural network parameterized with $\boldsymbol{w}$ and $\mathbf{K}_{\boldsymbol{\theta}}$ is the base kernel function (e.g., radial basis function) of GP. The deep learning transformation can capture the non-linear and hierarchical structure in high-dimensional data. The GP with the base kernel is applied on the final layer of DNN and make inference based on the learned latent features as Fig. 7 (a) shows. The idea has been successfully applied to spatio-temporal crop yield prediction where GP plays a role in accounting for the spatio-temporal auto-correlation between samples (You et al., 2017), which cannot be reflected from the DNN features.
@@
Use less dataset for training.
@@
Sample-density aware model Pros: Capture dataset shift uncertainty. Cons: Hard to ensure certain properties
@@
Distance-aware neural network: Though the modern neural networks have the capability of extracting representative features from large datasets, they are not aware of the extent of the distinction of new test samples to the training datasets. Thus, to characterize the uncertainty resulting from sample feature density, many approaches aim to take the distance awareness between samples into the design process of neural network motivated by Gaussian process (Liu et al., 2020). Assume the input data manifold is equipped with a metric $||\cdot||_{\mathcal{X}}$, which can quantify the distance between samples in the feature space. The intuition of a distance-aware neural network is to leverage the feature extraction capability of DNN to learn a hidden representation $h(\boldsymbol{x})$ that reflects a meaningful distance in the data manifold $||\boldsymbol{x}-\boldsymbol{x^{\prime}}||_{\mathcal{X}}$ as shown in Fig 8. However, one significant issue with the unconstrainted DNN model is the feature collapse, which means DNN feature extraction can map in-distribution data (training samples) and out-of-distribution data (lies further to the training data) to similar latent representations. Thus the Gaussian process based on the DNN extracted feature can be over-confident for those samples that lie further away from training samples.
@@
In summary, we summarize and compare existing works on uncertainty quantification arising from model uncertainty. The advantages and disadvantages of each method are concluded in Table 1. BNN model can capture model uncertainty arising from parameter estimation, but usually have high computational costs. The ensemble models can capture uncertainty from multiple perspectives, such as model architecture misspecification, low training dataset, and hyperparameters. The method also take high computational cost. On the other hand, the sample-density aware model can capture dataset shift uncertainty, but it’s often hard to learn the distance-aware feature space and need to add constraints to the neural network model.
@@ 5.2.1. Deep discriminative model
To quantify the data uncertainty, a discriminative model directly outputs a predictive distribution with a neural network. Specifically, the distribution can be represented by a parametric or non-parametric model. The parametric model assumes the output has an explicit parameterized family of probability distributions whose parameters (e.g., mean and variance for Gaussian distribution) are predicted with the neural network while the non-parametric model does not have any assumption on the underlying distributions. We will discuss existing works for each category in detail.
Parametric model: The standard approach for quantifying data uncertainty is directly learning a parametric model for $p(y|\boldsymbol{x},\boldsymbol{\theta})$. From a frequentist view, there exist a single set of optimum parameters $*\boldsymbol{\theta}*$. For the classification problem, $p(y|\boldsymbol{x},\boldsymbol{\theta})$ is a parameterized categorical distribution on the $k$ class, and the distribution parameter $\boldsymbol{\pi}=(\pi_{1},...,\pi_{K})$ are predicted from model output as Eq. 14 shows.
@@
For regression problems, data uncertainty is assumed to come from the inherent noise (or measurement error, human labeling error) in the training data. In general, the training data is modeled as independent additive Gaussian noise with sample-dependent variance $\sigma(\boldsymbol{x})$, which indicates the target variable $y_{i}=f_{\boldsymbol{\theta}}(\boldsymbol{x}_{i})+\epsilon(\boldsymbol{x}_{i})$. $\epsilon(\boldsymbol{x}_{i})$ is the independent heterogeneous Gaussian noise, which represents the uncertainty for each sample. In this way, the output will be a parameterized continuous Gaussian distribution, as Eq. 15 and Fig. 9 (a) shows. Both the mean and variance are predicted from the neural network (Kendall and Gal, 2017). The mean represents the prediction of the model and the variance represents the uncertainty of the sample prediction. To optimize the neural network parameters $\boldsymbol{\theta}$, maximize likelihood is performed on the mean and variance jointly as Eq. 15 shows. This is also known as heteroscedastic regression, which assumes the observational noise level varies with different samples. This is suitable for the case where some samples have higher noise (uncertainty), while others have lower. Besides Gaussian distribution, the neural network can be parameterized with many other kinds of distributions, such as mixture Gaussian distribution (Guillaumes, 2017), which is implemented with mixture density network (MDN) (Bishop, 1994), assuming multiple modes for the prediction. MDN has the advantage that can account for the uncertainty from multiple prediction modes but consumes more computation. It’s important to choose a suitable parameterized distribution, which depends on the nature of the problem.
@@
(15) $\footnotesize\begin{split}&p(y|\boldsymbol{x},\boldsymbol{\theta})=\mathcal{N}(f_{\boldsymbol{\theta}}(\boldsymbol{x}),\sigma_{\boldsymbol{\theta}}(\boldsymbol{x}))\ \ \text{and}\ \ \mathcal{L}_{\text{NN}}(\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2\sigma_{\boldsymbol{\theta}}(\boldsymbol{x}_{i})}^{2}||y_{i}-f_{\boldsymbol{\theta}}(\boldsymbol{x_{i}})||^{2}+\frac{1}{2}\log\sigma_{\boldsymbol{\theta}}(\boldsymbol{x}_{i})^{2}\end{split}$
@@
The advantage of prediction distribution is that it is simple to add the approach to existing neural network architecture and requires little modification to the training and inference process. However, the explicit parameterization form requires choosing the appropriate distribution to accurately capture the underlying uncertainty, which can be hard if no any prior information available.
@@
Non-parametric model: Another widely popular approach to indicate the data uncertainty is through prediction interval (PI) (Pearce et al., 2018; Wu et al., 2021). For regression problems, the prediction intervals output a lower and upper bound $[y_{l},y_{u}]$, with which we expect the ground truth $y$ falls into the interval with a prescribed confidence level, $1-\alpha$, meaning that $p(y\in[y_{l},y_{u}])>1-\alpha$ as shown in Fig. 9 (b). This approach does not require explicit distribution over the prediction variable and is more flexible. Traditional prediction intervals are constructed in two steps, First step is to learn the point estimation of the target variable, which is obtained through the minimization of the error-based loss function (i.e., mean square loss), and then estimate the prediction variance around the local optimum prediction. The strategy is trying to minimize the prediction error, but not optimize the prediction interval quality. One recent paper explicitly constructs a lower and upper bound estimation (LUBE) to improve the PI characteristics, i.e., the width and coverage probability. The basic intuition is that the PI should cover the ground-truth with a certain pre-defined probability (confidence level), but should be as narrow as possible. The approach improves the quality of the constructed PI, but the new cost function is non-differentiable and requires Simulated Annealing (SA) sampling to obtain the optimal NN parameters.
@@
To overcome the non-differentiable limitation of LUBE, another approach propose a coverage width-based loss function (Pearce et al., 2018) with a similar goal as LUBE, shown in the Eq.16. The mean prediction interval width (MPIW) is equal to $|y_{u}-y_{l}|$, and prediction interval coverage width (PICP) indicates the average probability that the PI covers the ground truth. The total loss encourages the PI to be narrow while having a higher coverage probability than the prescribed confidence level $\alpha$.
@@
Recent approaches put forward the method of prediction interval learning as an empirical constrained optimization problem. There are two different views of this optimization problem: primal and dual perspectives. The primal perspective views the optimization objective as minimizing the PI intervals under the constraints that the PI attains a coverage probability larger than the confidence level (Chen et al., 2021b), which is expressed as following:
@@
They presented a discriminative learning framework that optimizes the expected error rate under a budget constraint on the interval sizes. In this way, the approach avoids single-point loss and can provide a statistical guarantee of the generalization capability on the whole population. Compared with the primal setup, the dual perspective in a batch learning can construct the prediction interval of a group of test points at the same time and avoid the computational overhead.

### v5

@@ 5.1.2. Ensemble models
Ensemble models combine multiple neural network models in the prediction process. The ensemble predictions form an output distribution. The prediction variability of the ensemble models can be a measure of model uncertainty (e.g., a larger variance implies larger uncertainty). To capture the model uncertainty arising from different aspects, several strategies for constructing ensembles can be adopted, summarized into three major categories: the first kind is through bootstrapping (Lakshminarayanan et al., 2017). This is a strategy for random sampling from the original dataset with replacement. An ensemble of neural network models is constructed, each trained on different bootstrapped samples. After training, the inference is done by aggregating the ensembles, and uncertainty is obtained from the prediction variance (regression) or average entropy (classification). The second strategy is to construct different neural network architectures (number of layers, hidden neurons, type of activation functions.) (Mallick et al., 2022). This strategy can account for the uncertainty from model misspecification. Other strategies include different initialization of parameters along with a random shuffle of the datasets. This is better than the bootstrap strategy since more samples are utilized for each model. The third type is hyperensemble (Wenzel et al., 2020). This approach constructs ensembles with different hyper-parameters, such as learning rate, optimization strategy, and training strategy.
@@
5.1.3. Sample distribution-related methods
In Section 3.1.1, we described model uncertainty related to sample distribution as a third category. It further includes two cases: (1) test samples and training samples follow different distributions (out of distribution); (2) a test sample is far away from other training samples (or is surrounded by sparse training samples) in the feature space. Here we only review methods in the case (2), i.e., sample density (distance)-aware neural networks. We do not specifically discuss out-of-distribution (OOD) methods as BNNs and ensemble methods can also be used for OOD-related uncertainty. We will review the methods for OOD problems in Section 6.
Because of the theoretical soundness of the BNN model and the simplicity of deep ensemble models, they are the popular methods for modern DNN uncertainty quantification. However, the computational challenge with BNN and ensemble models make it infeasible for real-world applications. In addition, the approaches can address the model uncertainty associated with the parameter, model architecture, and training process stochasticity, but cannot generalize to model uncertainty coming from low sample density, which means, the samples lie far away from the support of the training sets may show over-confident results. In this regard, many approaches have been motivated to develop sample density-aware neural networks to capture the model uncertainty due to low training sample density. Existing work can be summarized into two categories: Gaussian process hybrid neural network and distance-aware neural network. Each category will be discussed in the following.
@@ 5.1.3. Sample distribution-related methods
Then the prediction on new test input sample $\boldsymbol{x}_{*}$ is through computing the posterior distribution conditioned on the training data $\mathcal{D}_{\text{tr}}$, given by Eq. 10. GP inference gives lower uncertainty if the test samples are around the region where training samples are abundant (higher sample density as in the middle part of Fig. 7), otherwise resulting in higher uncertainty (boundary part of Fig. 7). Note that even Gaussian Process belongs to Bayesian methods, it considers the uncertainty source different from the BNN methods, For the Gaussian process method, it captures the uncertainty related to sample sparsity. Thus, we categorize GP as a different submodule to BNN methods.
(10) $\footnotesize p(y_{*}|x_{*},\mathcal{D}_{\text{train}},\boldsymbol{\theta})=\mathcal{N}(y|\mathbf{K}_{\boldsymbol{x}}^{T}\mathbf{K}_{n}^{-1}\boldsymbol{y},K_{x*}-\mathbf{K}_{\boldsymbol{x}}^{T}\mathbf{K}_{n}^{-1}\mathbf{K}_{\boldsymbol{x}}),$
@@
Sparse Gaussian process: Though possessing the intriguing property of inherent uncertainty estimation, GP is prohibitive for large datasets because the inversion of covariance matrix requires $\mathcal{O}(n^{3})$ complexity ($n$ is the total number of training samples), which takes high computation and storage cost for training and inference on large datasets. In this regard, many methods (Snelson and Ghahramani, 2005; Titsias, 2009)attempt to make a sparse approximation to the full GP to bring down the computation to $\mathcal{O}(m^{2}n)$ ($m$ is the number of inducing variables and $m\ll n$. The inducing variables can be anywhere in the input domain, not constrained to be a subset of the training data, which is denoted as input-output pairs $\{\hat{\boldsymbol{x}}_{i},\hat{y}_{i}\}_{i=1}^{m}$. Then the inversion of the original covariance matrix $\mathbf{K}_{n}$ can be replaced with a low-rank approximation from the inducing variables, which only requires the inversion of $m\times m$ matrix $\mathbf{K}_{m}$. Then the question becomes how to select $m$ best-inducing variables to be representative of the training dataset. Common approaches assume that the best representative inducing variables are those that maximize the likelihood (ML) of the training data (Snelson and Ghahramani, 2005). Then the location of inducing variables and hyper-parameters of GP are optimized simultaneously through ML. The training data likelihood can be obtained by marginalizing the inducing variables on the joint distribution of training data and inducing variables. Another approach variation sparse GP formulates a variation lower bound of the exact likelihood by treating the locations of inducing variables as variational parameters (Titsias, 2009) for optimization. To further reduce the computation for more scalable inference on a large dataset, the paper (Wilson and Nickisch, 2015) proposes a kernel interpolation method for scalable structured GP method by exploiting the structure of the covariance by imposing grid constraint on the inducing variables. In this way, the kernel matrix $\mathbf{K}_{m}$ admits the Kronecker structure and is much easier for computing the inversion.
@@
Deep Gaussian process: Besides the computational challenge, another limitation of GP is the joint Gaussian distribution assumption on the target variables limits the model’s capability in capturing diverse relationships among instances in large datasets. Additionally, GP relies heavily on the kernel function to compute the similarity between samples by transforming input features into the high-dimensional manifold. However, for high-dimension structured data, it is challenging to construct appropriate kernel functions to extract hierarchical features for computing similarity between samples. To address these limitations, two lines of research areas have been proposed: deep kernel learning and Deep Gaussian process.
@@
Deep kernel learning (Wilson et al., 2016) aims to combine the structured feature learning capability of DNN with the GP to learn more flexible representations. The motivation is that DNN can automatically discover meaningful representations from high-dimensional data, which could alleviate the fixed kernel limitations of GP and improve its expressiveness. Specifically, the deep kernel learning transforms the kernel $\mathbf{K}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ to $\mathbf{K}_{\boldsymbol{\theta}}(g(\boldsymbol{x}_{i};\boldsymbol{w}),g(\boldsymbol{x}_{j},\boldsymbol{w}))$, where $g(\cdot;\boldsymbol{w})$ is the neural network parameterized with $\boldsymbol{w}$ and $\mathbf{K}_{\boldsymbol{\theta}}$ is the base kernel function (e.g., radial basis function) of GP. The deep learning transformation can capture the non-linear and hierarchical structure in high-dimensional data. The GP with the base kernel is applied on the final layer of DNN and makes inference based on the learned latent features as Fig. 8 (a) shows. The idea has been successfully applied to spatio-temporal crop yield prediction where GP plays a role in accounting for the spatio-temporal auto-correlation between samples (You et al., 2017), which cannot be reflected from the DNN features.
@@
Use fewer datasets for training.
@@
Sample distribution Pros: Capture dataset shift uncertainty. Cons: Hard to ensure certain properties
@@
Distance-aware neural network: Though modern neural networks have the capability of extracting representative features from large datasets, they are not aware of the extent of the distinction of new test samples to the training datasets. Thus, to characterize the uncertainty resulting from sample feature density, many approaches aim to take the distance awareness between samples into the design process of neural network motivated by Gaussian process (Liu et al., 2020). Assume the input data manifold is equipped with a metric $||\cdot||_{\mathcal{X}}$, which can quantify the distance between samples in the feature space. The intuition of a distance-aware neural network is to leverage the feature extraction capability of DNN to learn a hidden representation $h(\boldsymbol{x})$ that reflects a meaningful distance in the data manifold $||\boldsymbol{x}-\boldsymbol{x^{\prime}}||_{\mathcal{X}}$ as shown in Fig 9. However, one significant issue with the unconstrainted DNN model is the feature collapse, which means DNN feature extraction can map in-distribution data (training samples) and out-of-distribution data (lies further to the training data) to similar latent representations. Thus the Gaussian process based on the DNN extracted feature can be over-confident for those samples that lie further away from training samples.
@@
5.1.4. Summary on Model Uncertainty:
We summarize and compare existing methods for quantifying uncertainty arising from model uncertainty. The advantages and disadvantages of each method are concluded in Table 2. BNN model can capture model uncertainty arising from parameter estimation but usually has high computational costs. The ensemble models can capture uncertainty from multiple perspectives, such as model architecture misspecification, low training dataset, and hyperparameters. The method also has a high computational cost. On the other hand, the sample distribution-based model can capture dataset shift uncertainty, but it’s often hard to learn the distance-aware feature space and need to add constraints to the neural network model.
@@ 5.2.1. Deep discriminative model
To quantify the data uncertainty, a discriminative model directly outputs a predictive distribution with a neural network. Specifically, the distribution can be represented by a parametric or non-parametric model. The parametric model assumes the output follows an explicit parameterized family of probability distributions whose parameters (e.g., mean and variance for Gaussian distribution) are predicted with the neural network while the non-parametric model does not have any assumption on the underlying distributions. We will discuss existing works for each category in detail.
Parametric model: The standard approach for quantifying data uncertainty is directly learning a parametric model for $p(y|\boldsymbol{x},\boldsymbol{\theta})$. From a frequentist view, there exists a single set of optimum parameters $\boldsymbol{\theta}*$. For the classification problem, $p(y|\boldsymbol{x},\boldsymbol{\theta})$ is a parameterized categorical distribution on the $k$ class, and the distribution parameter $\boldsymbol{\pi}=(\pi_{1},...,\pi_{K})$ are predicted from model output as Eq. 13 shows.
@@
For regression problems, data uncertainty is assumed to come from the inherent noise (or measurement error, human labeling error) in the training data. In general, the training data is modeled as independent additive Gaussian noise with sample-dependent variance $\sigma(\boldsymbol{x})$, which indicates the target variable $y_{i}=f_{\boldsymbol{\theta}}(\boldsymbol{x}_{i})+\epsilon(\boldsymbol{x}_{i})$. $\epsilon(\boldsymbol{x}_{i})$ is the independent heterogeneous Gaussian noise, which represents the uncertainty for each sample. In this way, the output will be a parameterized continuous Gaussian distribution, as Eq. 14 and Fig. 10 (a) show. Both the mean and variance are predicted from the neural network (Kendall and Gal, 2017). The mean represents the prediction of the model and the variance represents the uncertainty of the sample prediction. To optimize the neural network parameters $\boldsymbol{\theta}$, maximize likelihood is performed on the mean and variance jointly as Eq. 14 shows. This is also known as heteroscedastic regression, which assumes the observational noise level varies with different samples. This is suitable for the case where some samples have higher noise (uncertainty), while others have lower. Besides Gaussian distribution, the neural network can be parameterized with many other kinds of distributions, such as mixture Gaussian distribution (Guillaumes, 2017), which is implemented with mixture density network (MDN) (Bishop, 1994), assuming multiple modes for the prediction. MDN has the advantage that can account for the uncertainty from multiple prediction modes but consumes more computation. It’s important to choose a suitable parameterized distribution, which depends on the nature of the problem.
@@
(14) $\footnotesize\begin{split}&p(y|\boldsymbol{x},\boldsymbol{\theta})=\mathcal{N}(f_{\boldsymbol{\theta}}(\boldsymbol{x}),\sigma_{\boldsymbol{\theta}}(\boldsymbol{x}))\ \ \text{and}\ \ \mathcal{L}_{\text{NN}}(\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2\sigma_{\boldsymbol{\theta}}^{2}(\boldsymbol{x}_{i})}||y_{i}-f_{\boldsymbol{\theta}}(\boldsymbol{x_{i}})||^{2}+\frac{1}{2}\log\sigma_{\boldsymbol{\theta}}^{2}(\boldsymbol{x}_{i})\end{split}$
@@
The advantage of prediction distribution is that it is simple to add the approach to existing neural network architecture and requires little modification to the training and inference process. However, the explicit parameterization form requires choosing the appropriate distribution to accurately capture the underlying uncertainty, which can be hard if no prior information is available.
@@
Non-parametric model: Another widely popular approach to indicate the data uncertainty is through prediction interval (PI) (Pearce et al., 2018; Wu et al., 2021). For regression problems, the prediction intervals output a lower and upper bound $[y_{l},y_{u}]$, with which we expect the ground truth $y$ falls into the interval with a prescribed confidence level, $1-\alpha$, meaning that $p(y\in[y_{l},y_{u}])>1-\alpha$ as shown in Fig. 10 (b). This approach does not require explicit distribution over the prediction variable and is more flexible. Traditional prediction intervals are constructed in two steps, First step is to learn the point estimation of the target variable, which is obtained through the minimization of the error-based loss function (i.e., mean square loss), and then estimate the prediction variance around the local optimum prediction. The strategy is trying to minimize the prediction error, but not optimize the prediction interval quality. One recent paper explicitly constructs a lower and upper bound estimation (LUBE) to improve the PI characteristics, i.e., the width and coverage probability. The basic intuition is that the PI should cover the ground truth with a certain pre-defined probability (confidence level), but should be as narrow as possible. The approach improves the quality of the constructed PI, but the new cost function is non-differentiable and requires Simulated Annealing (SA) sampling to obtain the optimal NN parameters.
@@
To overcome the non-differentiable limitation of LUBE, another approach proposes a coverage width-based loss function (Pearce et al., 2018) with a similar goal as LUBE, shown in the Eq.15. The mean prediction interval width (MPIW) is equal to $|y_{u}-y_{l}|$, and prediction interval coverage width (PICP) indicates the average probability that the PI covers the ground truth. The total loss encourages the PI to be narrow while having a higher coverage probability than the prescribed confidence level $\alpha$.
@@
Recent approaches put forward the method of prediction interval learning as an empirical constrained optimization problem. There are two different views of this optimization problem: primal and dual perspectives. The primal perspective views the optimization objective as minimizing the PI intervals under the constraints that the PI attains a coverage probability larger than the confidence level (Chen et al., 2021b), which is expressed as follows:
@@
They presented a discriminative learning framework that optimizes the expected error rate under a budget constraint on the interval sizes. In this way, the approach avoids single-point loss and can provide a statistical guarantee of the generalization capability of the whole population. Compared with the primal setup, the dual perspective in batch learning can construct the prediction interval of a group of test points at the same time and avoid the computational overhead.